HETAV PATEL

Grade 11,

A weak acid HA after treatment with 12 ml of 0.1 M stong base BOH has a pH of 5.At the end point,the volume of same base required is 26.6 ml.Calculate K a of acid.

A weak acid HA  after treatment with 12 ml of 0.1 M stong base BOH has a pH of 5.At the end point,the volume of same base required is 26.6 ml.Calculate Ka of acid.

Grade:12th Pass

1 Answers

User
5 years ago
Total milliequivalents of acid = milliequivalents of base = 26.6 x 0.1 = 2.66But the resultant mixture has HA leftThe reaction is as follows: HA + BOH \rightarrow BA + H2Ot=0 2.66 12 0 0t1 1.46 0 1.2 1.2Therefore for a buffer solution,pH = pKa+\log \frac{\left [ salt \right ]}{\left [ acid \right ]}\Rightarrow 5 = pKa+\log \frac{1.2}{1.46}\Rightarrow \log Ka = -5.86\Rightarrow Ka = 8.21 \times 10^{-6}